What is the sum of three consecutive odd integers if the sum of the second and third numbers is six more than two times the first number?
Answer by algebrahouse.com(1659)

"What is the sum of three consecutive odd integers if the sum of the second and third numbers is six more than two times the first number?"

x = first odd integer
x + 2 = 2nd odd integer
x + 4 = 3rd odd integer

x + 2 + x + 4 = 2x + 6 {sum of 2nd and 3rd is 6 more than two times first}
2x + 6 = 2x + 6 {combined like terms}
6 = 6 {subtracted 2x from both sides}
= infinite solutions

*When the variable disappears and you are left with a true statement,
there are infinite solutions for x*

In other words, x could be any odd number.

The statement "the sum of the second and third numbers is six more than two times the first number" is true for any three consecutive odd integers.


For more help from me, visit: www.algebrahouse.com